# Eccentricity distribution of extreme mass ratio inspirals

### v1

@@ II.2 Astrophysical formation rates
We employ loss cone theory to weight EMRIs based on their semi-major axis $a$ at the start of the runs performed in Mancieri et al. [67] (see Fig. 1), depending on their formation rate in the nuclear star cluster. As in Mancieri et al. [67], we solve the orbit averaged Fokker-Planck equation to compute the time evolution of the distribution function of a spherical stellar system because of two-body relaxation [83].
Loss cone theory provides a framework to compute the rate of particles pushed at pericenter smaller than the loss cone radius $R_{\mathrm{lc}}=8\,R_{\mathrm{g}}$. Considering that the Fokker-Planck coefficients are computed in Newtonian dynamics, the value of $R_{\mathrm{lc}}$ is chosen so that a highly eccentric orbit with such a pericenter in the potential $V_{\mathrm{Newt}}(r)=-GM_{\bullet}/r+L^{2}/2r^{2}$ has $L=4\,c\,R_{\mathrm{g}}$, the limiting value prescribed by general relativity for a particle to avoid plunging on a Schwarzschild MBH [84, 85, 67].
Because of the stellar potential, we label the specific energy in the potential of the nuclear star cluster $\mathcal{E}$, and infer the semi-major axis $a(\mathcal{E})$ from the radius of the circular orbit $r_{\mathrm{c}}$ at that energy, using the same mapping as in Mancieri et al. [67]. The rate $\mathrm{d}\dot{N}_{\mathrm{lc}}$ of objects captured with energy between $\mathcal{E}$ and $\mathcal{E}+\mathrm{d}\mathcal{E}$ can be expressed in terms of the differential rate $\mathcal{F}_{\mathrm{lc}}$, also known as the loss cone flux:
@@
$S(\mathcal{E})$ is always $1$ for $a\to 0$; at larger $a$ it decreases to a lower value, which is zero for $M_{\bullet}\gtrsim 3\times 10^{5}\,\mathrm{M_{\odot}}$ [91, 92], but larger than zero for lower central MBH masses [93, 67]. In particular, we use the numerical $S(\mathcal{E})$ presented in Mancieri et al. [67] [^4] See Sect. 4.4 and specifically Eq. 61 in Mancieri et al. [67]..
We simulate the relaxation process of stars and stellar-mass BHs within nuclear star clusters surrounding MBHs with mass $M_{\bullet}\in$$10^{4}$, $10^{5}$, $3\times 10^{5}$, $10^{6}$, $4\times 10^{6}$} $\mathrm{M}_{\odot}$ with the software presented in Broggi et al. [86]. The stellar distribution is characterized by two components: stars of $1\,\mathrm{M}_{\odot}$ each are contributing to a total mass of $20M_{\bullet}$, while stellar-mass BHs of $10\,\mathrm{M}_{\odot}$ each contribute in total to $0.2M_{\bullet}$. We initialize both components on the same Dehnen profile [94, 95] with inner slope $1.5$ and scale radius $4\,R_{\mathrm{inf}}$. The initial slopes align with values commonly employed in literature (e.g. Refs. [96, 86]), and are consistently evolved with time following mass segregation. The scale radius of the distribution is chosen such that the stellar cluster surrounding our $4\times 10^{6}\,\mathrm{M_{\odot}}$ matches the observed density at the influence radius of SgrA (e.g. Refs [8]). For each system, we compute $\mathcal{F}_{\mathrm{EMRI}}$ as a function of time and stop the simulation at $t_{\mathrm{end}}$, which is $10$ times the time-to-peak $t_{\mathrm{peak}}$ of the EMRI rate [^5] In the case of $M_{\bullet}=4\times 10^{6}\,\mathrm{M_{\odot}}$, we stop at $t_{\mathrm{end}}=10\,\mathrm{Gyr}$, since $t_{\mathrm{peak}}\approx 2\,\mathrm{Gyr}$.. Indeed, as $\mathcal{F}_{\mathrm{lc}}$ depends on time, so does $\mathcal{F}_{\mathrm{EMRI}}$: Broggi et al. [86] shows that the EMRI rates reach maxima (which are in general agreement with literature based on a steady-state assumption) at $t_{\mathrm{peak}}$ and then undergo a nearly exponential decay.
We introduce a log uniform distribution in semi-major axis centered at the values $a_{\mathrm{i}}$ used as initial conditions in Mancieri et al. [67]. This grid in $a$ is then mapped to energy space using the energy of the corresponding circular orbit. We compute the average rate of EMRIs produced in the $i$-th energy cell as
@@
In Fig. 3 we consider the case $M_{\bullet}=10^{5}\,M_{\odot}$ and compare the fraction of EMRIs in the sample that form with semi-major axis $a_{i}$ at time $t=\{1/3,\ 1,\ 3,\ 10\}\,t_{\mathrm{peak}}$, together with the time average. At later times larger values of $a_{i}$ are increasingly favored. In fact, after the initial migration of BHs towards the central region because of mass segregation, at time $t\simeq t_{\mathrm{peak}}$ the system enters a phase of expansion driven by diffusion. This shifts the the peak of $\mathcal{F}_{\mathrm{lc}}(t,\mathcal{E})$ to orbits with larger and larger semi-major axes as relaxation proceeds [86].
[…]
@@ III Results: EMRI eccentricity distributions
Figure 4: EMRI eccentricity distribution at plunge, for different central MBH masses. Here we report the overall probability density function of our EMRI population (blue histograms) and display the underneath distributions of “classical” and “initially wide” EMRIs (green and red shaded histograms, see Sect. III for details), which are normalized to recover the overall distribution when summed together. We also show (orange dashed histogram) the distribution reported in Babak et al. [27], which has become a standard assumption in the literature.
We also separately display the distributions for “classical EMRIs” and “initially wide EMRIs”. We make this distinction based on the initial $a_{\mathrm{i}}$ value used to initialize runs in Mancieri et al. [67]. We classify the former (latter) as those such that $a_{\mathrm{i}}<a_{\mathrm{c}}$ ($a_{\mathrm{i}}>a_{\mathrm{c}}$). Here $a_{\mathrm{c}}$ is the critical semi-major axis at which the probability of forming EMRIs drops from 1 (at small $a$) to its value at large $a$. We define $a_{\mathrm{c}}$ such that
@@
using the numerical $S(a)$ found in Mancieri et al. [67] for each $M_{\bullet}$ case. The approximate values for $a_{\mathrm{c}}$ are $5\times 10^{-3},9\times 10^{-3},10^{-2},10^{-2},10^{-2}$} $R_{\mathrm{inf}}$ for $M_{\bullet}=$$10^{4}$, $10^{5}$, $3\times 10^{5}$, $10^{6}$, $4\times 10^{6}$} $\mathrm{M}_{\odot}$ (see Fig. 1). We make this distinction because classically EMRIs are thought to only form for $a$ strictly less than $a_{\mathrm{c}}$, while a more thorough analysis shows that there is a smooth transition from the values of $a$ which allow for EMRI formation and those that do not, encompassing about an order of magnitude in $a$ [67]. The wide EMRIs sub-population also includes (and for $M_{\bullet}\leq 3\times 10^{5}\,\mathrm{M_{\odot}}$ is mostly represented by) “cliffhanger EMRIs” [93, 67]. As the flux of objects crossing into the loss cone typically peaks at $a>a_{\mathrm{c}}$ [86], these initially wide EMRIs can represent a significant fraction of the total. The only case where the flux peaks at $a<a_{\mathrm{c}}$ is that of $M=10^{4}\,\mathrm{M_{\odot}}$. This is likely the reason why for this case only we observe a peak at lower eccentricity. In fact, the flux peaking at lower $a$ means that the binaries that are weighted more are those entering the GW-dominated region far away from the separatrix (see Fig. 1), and can thus circularize more before the final plunge. Figure 4 shows that the two populations of EMRIs have different eccentricity distributions, with wide EMRIs being more eccentric when they plunge, with peaks at $e_{\mathrm{pl}}$ roughly double that of classical EMRIs. Despite this, the underlying populations are not separated enough for the overall distributions to show bi-modality.
In Fig. 5, we show the location on the $(p,e)$ plane of EMRIs at 3 months, 6 months, 1 year and 2 years before the plunge [^7] The points are obtained by interpolating linearly between the two time-steps of the integration to plunge enclosing the target time.. In some cases (especially for low $M_{\bullet}$) this required integrating backwards in time the final conditions of Mancieri et al. [67], and this led us to exclude $M_{\bullet}=10^{4}\,\mathrm{M_{\odot}}$ from this analysis (see below). The sharp transition that can be noticed for $M_{\bullet}=10^{6}\,\mathrm{M_{\odot}}$ and $M_{\bullet}=4\times 10^{6}\,\mathrm{M_{\odot}}$ is due to binaries plunging outside the region where the KerrEccEqFlux trajectory module can be applied, as the PN5 module overestimates the plunge time given the same initial conditions. We observe similar discontinuities even when switching between modules at later stages, much within the region for which numerical fluxes are available. We stress that the time to plunge is not an intrinsic physical property of the system, and there is no discontinuity in the evolution of the orbital parameters themselves, as shown both in Fig. 2 and by the thin red lines in Fig. 5 itself.
@@
[…]
As mentioned above, in Fig. 5, we exclude the case $M_{\bullet}=10^{4}\,\mathrm{M_{\odot}}$ from the analysis as we find that a large fraction of systems spend much less than 2 years in the region of the orbital parameters where two-body relaxation is negligible (where $t_{\mathrm{GW}}<10^{-3}\,t_{\mathrm{rlx}}$ in Fig. 1). This prevents trusting backward time integration from plunge for 2 years or more, as a significant fraction of systems was influenced by relaxation at that time. For $M_{\bullet}=10^{5}\,\mathrm{M_{\odot}}$, 13% of binaries in our population stay in the region where $t_{\mathrm{GW}}<10^{-3}\,t_{\mathrm{rlx}}$ for less than 2 years and 3% of them for less than 1 year. Binaries that plunge this quickly have small initial pericenter, which translates to large $e_{0}$ given that all of our binaries are initialized on the $t_{\mathrm{GW}}=10^{-3}\,t_{\mathrm{rlx}}$ curve shown in Fig. 1. In particular, for $M_{\bullet}=10^{5}\,\mathrm{M_{\odot}}$ binaries that take 1-2 years to merge have $1-e_{0}\approx 4\times 10^{-4}$, so that the distributions shown in Figs. 5 and 6 might be increasingly less accurate for eccentricities larger than such value; we nevertheless expect two-body relaxation to still be subdominant compared to GWs emission. For $M_{\bullet}\geq 3\times 10^{5}\,\mathrm{M_{\odot}}$, there is no such issue.
We also check if our binaries are within the LISA band by estimating their peak harmonic GW frequency $f_{\mathrm{GW}}^{\mathrm{peak}}$ during their inspiral, using the numerical fit from Hamers [97]:
@@ IV Comparison with previous works
The first estimate of EMRI eccentricity can be found in Hopman and Alexander [91], who reported the eccentricity distribution for the case of $M_{\bullet}=3\times 10^{6}\,\mathrm{M_{\odot}}$ at semi-major axis $a_{\mathrm{HA}}=3.24\times 10^{-6}\,\mathrm{pc}$ (about $22.6\,R_{\mathrm{g}}$), so that the Keplerian period is $10^{4}\,\mathrm{s}$, a threshold they used to estimate the entry in the LISA band. They performed Monte Carlo simulations, accounting for the effect of two-body relaxation and GWs emission on the energy and angular momentum of the orbits. The distribution they found is reported in Fig. 7. Despite not having performed simulations for $M_{\bullet}=3\times 10^{6}\,\mathrm{M_{\odot}}$, we reproduce with large agreement their results, using our data and weights for the $M_{\bullet}=4\times 10^{6}\,\mathrm{M_{\odot}}$ case and evolving until the same semi-major axis [^8] The eccentricity values are obtained by interpolating linearly between the two time-steps of the integration to plunge enclosing the condition $a(t)=a_{\mathrm{HA}}$. $a_{\mathrm{HA}}$. Our main results extend their work to the plunge, and for a larger range of MBH masses, crucially including the contribution of wide and cliffhanger EMRIs.
@@
Figure 7: EMRI eccentricity distribution for $M_{\bullet}=4\times 10^{6}\,\mathrm{M_{\odot}}$ at $a=a_{\mathrm{HA}}$, compared to the results by Hopman and Alexander [91] with $M_{\bullet}=3\times 10^{6}\,\mathrm{M_{\odot}}$ at the same semi-major axis. A small fraction ($0.14\%$) of EMRIs have already plunged before reaching $a_{\mathrm{HA}}$, and thus are not shown in this figure. They live in the small tail at $e_{\mathrm{pl}}\gtrsim 0.8$ in Fig. 4.
Distributions of $e_{\mathrm{pl}}$ spanning a range of MBH masses have been recently presented by Rom et al. [98]. They first derive a steady-state stellar distribution surrounding the MBH by analytical means, assuming a population of stars of mass $m_{\star}=1\,\mathrm{M_{\odot}}$, and one of stellar-mass BHs of mass $m_{\bullet}=10\,\mathrm{M_{\odot}}$. In their work, stars and BHs are on profiles with inner slopes $1.5$ and $1.75$ respectively in the region of EMRI formation. In our simulations, the initial inner slope for both is $1.5$, but at the final stage the two components do not fit a power law profile. Nonetheless, at the mpc scale they are locally described by inner slope $1.3$ for stars and $1.7$ for BHs [86].
@@
For $f_{\bullet}>f_{\mathrm{c}}$, they report that the distribution stabilizes on [^10] The result was first derived by Linial and Sari [99] in the context of stellar EMRIs. $P(e_{\mathrm{pl}})=2g(e_{\mathrm{pl}})g^{\prime}(e_{\mathrm{pl}})$, now being independent on the value of $f_{\bullet}$, where
@@
and $g^{\prime}(e)$ is its derivative with respect to $e$. Here they make use of [100]
@@
Although the general trends of their distribution are qualitatively comparable to ours, they feature a peak at smaller $e_{\mathrm{pl}}$ and a slower decay as $e_{\mathrm{pl}}\rightarrow 1$ compared to what we observe. Even for $M_{\bullet}=10^{4}\,\mathrm{M_{\odot}}$, for which the location of the peaks is compatible, our decay at large eccentricity is better fitted by an exponential as opposed to a power law. The incompatibility is also caused by the fact that their $P(e_{\mathrm{pl}})$ does not smoothly go to 0 when $e_{\mathrm{pl}}\to 1$, and instead is sharply cut at some $e_{\mathrm{max}}$. For example, their model predicts that (independently on the MBH mass) $1\%$ of EMRIs plunge at $e_{\mathrm{pl}}>0.9$ if $f_{\bullet}=10^{-4}$, and $7\%$ do the same if $f_{\bullet}>f_{\mathrm{c}}$. For comparison, we get between $0.1\%$ and $0.5\%$ depending on $M_{\bullet}$ for the same computation.

### v2

@@ II.2 Astrophysical formation rates
[…]
Figure 3: In the upper panel, time dependent density profiles for stars (blue) and stellar-mass BHs (orange) in our nuclear star cluster model with an MBH of $10^{5}\,\mathrm{M_{\odot}}$. Both components are initially distributed with an inner slope of 1.5 but different normalizations. Profiles are shown at three different times: the initial condition (dotted lines), $t_{\mathrm{peak}}$ (dashed lines), and $10\,t_{\mathrm{peak}}$ (solid lines). In the lower panel, time evolution of the local slope of the density profiles as a function of radius $r$.
We employ loss cone theory to weight EMRIs based on their semi-major axis $a$ at the start of the runs performed in 2025A&A...694A.272M (see Fig. 1), depending on their formation rate in the nuclear star cluster. As in 2025A&A...694A.272M, we solve the orbit averaged Fokker-Planck equation to compute the time evolution of the distribution function of a spherical stellar system because of two-body relaxation [1978ApJ...226.1087C].
Loss cone theory provides a framework to compute the rate of particles pushed at pericenter smaller than the loss cone radius $R_{\mathrm{lc}}=8\,R_{\mathrm{g}}$. Considering that the Fokker-Planck coefficients are computed in Newtonian dynamics, the value of $R_{\mathrm{lc}}$ is chosen so that a highly eccentric orbit with such a pericenter in the potential $V_{\mathrm{Newt}}(r)=-GM_{\bullet}/r+L^{2}/2r^{2}$ has $L=4\,c\,R_{\mathrm{g}}$, the limiting value prescribed by general relativity for a particle to avoid plunging on a Schwarzschild MBH [1994PhRvD..50.3816C, 2011PhRvD..84d4024M, 2025A&A...694A.272M].
Because of the stellar potential, we label the specific energy in the potential of the nuclear star cluster $\mathcal{E}$, and infer the semi-major axis $a(\mathcal{E})$ from the radius of the circular orbit $r_{\mathrm{c}}$ at that energy, using the same mapping as in 2025A&A...694A.272M. The rate $\mathrm{d}\dot{N}_{\mathrm{lc}}$ of objects captured with energy between $\mathcal{E}$ and $\mathcal{E}+\mathrm{d}\mathcal{E}$ can be expressed in terms of the differential rate $\mathcal{F}_{\mathrm{lc}}$, also known as the loss cone flux:
@@
$S(\mathcal{E})$ is always $1$ for $a\to 0$; at larger $a$ it decreases to a lower value, which is zero for $M_{\bullet}\gtrsim 3\times 10^{5}\,\mathrm{M_{\odot}}$ [2005ApJ...629..362H, 2021MNRAS.501.5012R], but larger than zero for lower central MBH masses [2024PhRvL.133n1401Q, 2025A&A...694A.272M]. In particular, we use the numerical $S(\mathcal{E})$ presented in 2025A&A...694A.272M [^4] See Sect. 4.4 and specifically Eq. 61 in 2025A&A...694A.272M..
We simulate the relaxation process of stars and stellar-mass BHs within nuclear star clusters surrounding MBHs with mass $M_{\bullet}\in$$10^{4}$, $10^{5}$, $3\times 10^{5}$, $10^{6}$, $4\times 10^{6}$} $\mathrm{M}_{\odot}$ with the software presented in 2022MNRAS.514.3270B. The stellar distribution is characterized by two components: (i) stars of $1\,\mathrm{M}_{\odot}$ each, for a total stellar mass of $20M_{\bullet}$; (ii) stellar-mass BHs of $10\,\mathrm{M}_{\odot}$ each, for a total mass in BHs of $0.2M_{\bullet}$. We initialize both components on the same Dehnen profile [1993MNRAS.265..250D, 1994AJ....107..634T] with inner slope $1.5$ and scale radius $4\,R_{\mathrm{inf}}$. The initial slopes align with values commonly employed in the literature (e.g. Refs. [1977ApJ...216..883B, 2022MNRAS.514.3270B]), and are consistently evolved with time following mass segregation. The scale radius of the distribution is chosen such that the stellar cluster surrounding our $4\times 10^{6}\,\mathrm{M_{\odot}}$ matches the observed density at the influence radius of SgrA (e.g. Ref. [2007A&A...469..125S]). For each system, we compute $\mathcal{F}_{\mathrm{EMRI}}$ as a function of time and stop the simulation at $t_{\mathrm{end}}$, which is $10$ times the time-to-peak $t_{\mathrm{peak}}$ of the EMRI rate [^5] In the case of $M_{\bullet}=4\times 10^{6}\,\mathrm{M_{\odot}}$, we stop at $t_{\mathrm{end}}=10\,\mathrm{Gyr}$, since $t_{\mathrm{peak}}\approx 2\,\mathrm{Gyr}$.. Indeed, as $\mathcal{F}_{\mathrm{lc}}$ depends on time, so does $\mathcal{F}_{\mathrm{EMRI}}$: 2022MNRAS.514.3270B showed that the EMRI rates reach a maximum at $t_{\mathrm{peak}}$ and then undergo a nearly exponential decay. Moreover, the maximum was found to be in agreement steady state EMRI rate values quoted in the literature.
Figure 3 shows the time evolution of the density profiles for stars and stellar-mass BHs. Due to mass segregation, stars are displaced toward larger radii, while stellar-mass BHs initially sink toward the central MBH, resulting in a peak of EMRI formation at $t_{\mathrm{peak}}$. At later times, as the cluster continues to relax and expand, the BHs are also redistributed to larger radii. Consequently, the stellar density profile becomes progressively shallower in the inner region, whereas the BH profile first steepens and then slowly flattens.
[…]
We introduce a log uniform distribution in semi-major axis centered at the values $a_{\mathrm{i}}$ used as initial conditions in 2025A&A...694A.272M. This grid in $a$ is then mapped to energy space using the energy of the corresponding circular orbit. We compute the average rate of EMRIs produced in the $i$-th energy cell as
@@
In Fig. 4 we consider the case $M_{\bullet}=10^{5}\,M_{\odot}$ and compare the fraction of EMRIs in the sample that form with semi-major axis $a_{i}$ at time $t=\{1/3,\ 1,\ 3,\ 10\}\,t_{\mathrm{peak}}$, together with the time average. At later times larger values of $a_{i}$ are increasingly favored. In fact, after the initial migration of BHs towards the central region because of mass segregation, at time $t\simeq t_{\mathrm{peak}}$ the system enters a phase of expansion driven by diffusion. This shifts the the peak of $\mathcal{F}_{\mathrm{lc}}(t,\mathcal{E})$ to orbits with larger and larger semi-major axes as relaxation proceeds [2022MNRAS.514.3270B].
@@ III Results: EMRI eccentricity distributions
Figure 5: EMRI eccentricity distribution at plunge, for different central MBH masses. Here we report the overall probability density function of our EMRI population (blue histograms) and display the underneath distributions of “classical” and “initially wide” EMRIs (green and red shaded histograms, see Sect. III for details), which are normalized to recover the overall distribution when summed together. We also show (orange dashed histogram) the distribution reported in 2017PhRvD..95j3012B, which has become a standard assumption in the literature.
We also separately display the distributions for “classical EMRIs” and “initially wide EMRIs”. We make this distinction based on the initial $a_{\mathrm{i}}$ value used to initialize runs in 2025A&A...694A.272M. We classify the former (latter) as those such that $a_{\mathrm{i}}<a_{\mathrm{c}}$ ($a_{\mathrm{i}}>a_{\mathrm{c}}$). Here $a_{\mathrm{c}}$ is the critical semi-major axis at which the probability of forming EMRIs drops from 1 (at small $a$) to its value at large $a$. We define $a_{\mathrm{c}}$ such that
@@
using the numerical $S(a)$ found in 2025A&A...694A.272M for each $M_{\bullet}$ case. The approximate values for $a_{\mathrm{c}}$ are $5\times 10^{-3},9\times 10^{-3},10^{-2},10^{-2},10^{-2}$} $R_{\mathrm{inf}}$ for $M_{\bullet}=$$10^{4}$, $10^{5}$, $3\times 10^{5}$, $10^{6}$, $4\times 10^{6}$} $\mathrm{M}_{\odot}$ (see Fig. 1). We make this distinction because classically EMRIs are thought to only form for $a$ strictly less than $a_{\mathrm{c}}$, while a more thorough analysis shows that there is a smooth transition from the values of $a$ which allow for EMRI formation and those that do not, encompassing about an order of magnitude in $a$ [2025A&A...694A.272M]. The wide EMRIs sub-population also includes (and for $M_{\bullet}\leq 3\times 10^{5}\,\mathrm{M_{\odot}}$ is mostly represented by) “cliffhanger EMRIs” [2024PhRvL.133n1401Q, 2025A&A...694A.272M]. As the flux of objects crossing into the loss cone typically peaks at $a>a_{\mathrm{c}}$ [2022MNRAS.514.3270B], these initially wide EMRIs can represent a significant fraction of the total. The only case where the flux peaks at $a<a_{\mathrm{c}}$ is that of $M=10^{4}\,\mathrm{M_{\odot}}$. This is likely the reason why for this case only we observe a peak at lower eccentricity. In fact, the flux peaking at lower $a$ means that the binaries that are weighted more are those entering the GW-dominated region far away from the separatrix (see Fig. 1), and can thus circularize more before the final plunge. Figure 5 shows that the two populations of EMRIs have different eccentricity distributions, with wide EMRIs being more eccentric when they plunge, with peaks at $e_{\mathrm{pl}}$ roughly double that of classical EMRIs. Despite this, the underlying populations are not separated enough for the overall distributions to show bi-modality.
In Fig. 6, we show the location on the $(p,e)$ plane of EMRIs at 3 months, 6 months, 1 year and 2 years before the plunge [^7] The points are obtained by interpolating linearly between the two time-steps of the integration to plunge enclosing the target time.. In some cases (especially for low $M_{\bullet}$) this required integrating backwards in time the final conditions of 2025A&A...694A.272M, and this led us to exclude $M_{\bullet}=10^{4}\,\mathrm{M_{\odot}}$ from this analysis (see below). The sharp transition that can be noticed for $M_{\bullet}=10^{6}\,\mathrm{M_{\odot}}$ and $M_{\bullet}=4\times 10^{6}\,\mathrm{M_{\odot}}$ is due to binaries plunging outside the region where the KerrEccEqFlux trajectory module can be applied, as the PN5 module overestimates the plunge time given the same initial conditions. We observe similar discontinuities even when switching between modules at later stages, much within the region for which numerical fluxes are available. We stress that the time to plunge is not an intrinsic physical property of the system, and there is no discontinuity in the evolution of the orbital parameters themselves, as shown both in Fig. 2 and by the thin red lines in Fig. 6 itself.
@@
As mentioned above, in Fig. 6, we exclude the case $M_{\bullet}=10^{4}\,\mathrm{M_{\odot}}$ from the analysis as we find that a large fraction of systems spend much less than 2 years in the region of the orbital parameters where two-body relaxation is negligible (where $t_{\mathrm{GW}}<10^{-3}\,t_{\mathrm{rlx}}$ in Fig. 1). This prevents trusting backward time integration from plunge for 2 years or more, as a significant fraction of systems was influenced by relaxation at that time. For $M_{\bullet}=10^{5}\,\mathrm{M_{\odot}}$, 13% of binaries in our population stay in the region where $t_{\mathrm{GW}}<10^{-3}\,t_{\mathrm{rlx}}$ for less than 2 years and 3% of them for less than 1 year. Binaries that plunge this quickly have small initial pericenter, which translates to large $e_{0}$ given that all of our binaries are initialized on the $t_{\mathrm{GW}}=10^{-3}\,t_{\mathrm{rlx}}$ curve shown in Fig. 1. In particular, for $M_{\bullet}=10^{5}\,\mathrm{M_{\odot}}$ binaries that take 1-2 years to merge have $1-e_{0}\approx 4\times 10^{-4}$, so that the distributions shown in Figs. 6 and 7 might be increasingly less accurate for eccentricities larger than such value; we nevertheless expect two-body relaxation to still be subdominant compared to GWs emission. For $M_{\bullet}\geq 3\times 10^{5}\,\mathrm{M_{\odot}}$, there is no such issue.
We also check if our binaries are within the LISA band by estimating their peak harmonic GW frequency $f_{\mathrm{GW}}^{\mathrm{peak}}$ during their inspiral, using the numerical fit from 2021RNAAS...5..275H:
@@ IV Comparison with previous works
The first estimate of EMRI eccentricity can be found in 2005ApJ...629..362H, who reported the eccentricity distribution for the case of $M_{\bullet}=3\times 10^{6}\,\mathrm{M_{\odot}}$ at semi-major axis $a_{\mathrm{HA}}=3.25\times 10^{-6}\,\mathrm{pc}$ (about $22.6\,R_{\mathrm{g}}$), so that the Keplerian period is $10^{4}\,\mathrm{s}$, a threshold they used to estimate the entry in the LISA band. They performed Monte Carlo simulations, accounting for the effect of two-body relaxation and GWs emission on the energy and angular momentum of the orbits. Given a semi-major axis $a$, the maximum eccentricity a binary might have without plunging is [2005ApJ...629..362H]
$e_{\mathrm{max}}(a)=-\frac{R_{\mathrm{g}}}{a}+\sqrt{1-\frac{6R_{\mathrm{g}}}{a}+\frac{R_{\mathrm{g}}^{2}}{a^{2}}}\,$ (11)
which follows from Eq. (2) and $p=a(1-e^{2})$. Consistently, their distribution (reported in Fig. 8) shows a maximum eccentricity around $e_{\mathrm{max}}(a_{\mathrm{HA}})=0.81$. Despite not having performed simulations for $M_{\bullet}=3\times 10^{6}\,\mathrm{M_{\odot}}$, we reproduce with large agreement their results, using our data and weights for the $M_{\bullet}=4\times 10^{6}\,\mathrm{M_{\odot}}$ case and evolving until the same semi-major axis [^8] The eccentricity values are obtained by interpolating linearly between the two time-steps of the integration to plunge enclosing the condition $a(t)=a_{\mathrm{HA}}$. $a_{\mathrm{HA}}$. Our main results extend their work to the plunge, and for a larger range of MBH masses, crucially including the contribution of wide and cliffhanger EMRIs.
@@
Figure 8: EMRI eccentricity distribution for $M_{\bullet}=4\times 10^{6}\,\mathrm{M_{\odot}}$ at $a=a_{\mathrm{HA}}$, compared to the results by 2005ApJ...629..362H with $M_{\bullet}=3\times 10^{6}\,\mathrm{M_{\odot}}$ at the same semi-major axis. A small fraction ($0.14\%$) of EMRIs have already plunged before reaching $a_{\mathrm{HA}}$, and thus are not shown in this figure. They live in the small tail at $e_{\mathrm{pl}}\gtrsim 0.8$ in Fig. 5.
2021MNRAS.501.5012R present eccentricity distributions for $M_{\bullet}=4\times 10^{6}\,\mathrm{M_{\odot}}$, following an approach similar to that of 2005ApJ...629..362H and finding similar results. However, they terminate their simulations when the orbital period reaches $10^{3}\,\mathrm{s}$, which corresponds to a final semi-major axis of $a_{\mathrm{RP}}=7.71\times 10^{-7}\,\mathrm{pc}$ (approximately $4.03\,R_{\mathrm{g}}$). Since this radius lies within the Schwarschild separatrix for all values of $e$, we cannot perform a direct comparison of the eccentricity distribution.
Distributions of $e_{\mathrm{pl}}$ spanning a range of MBH masses have been recently presented by 2024ApJ...977....7R. They first derive a steady-state stellar distribution surrounding the MBH by analytical means, assuming a population of stars of mass $m_{\star}=1\,\mathrm{M_{\odot}}$, and one of stellar-mass BHs of mass $m_{\bullet}=10\,\mathrm{M_{\odot}}$. In their work, stars and BHs are on profiles with inner slopes $1.5$ and $1.75$ respectively in the region of EMRI formation. In our simulations, the initial inner slope for both is $1.5$, but at the final stage the two components do not fit a power law profile. Nonetheless, at the mpc scale they are locally described by inner slope $1.3$ for stars and $1.7$ for BHs [2022MNRAS.514.3270B].
@@
For $f_{\bullet}>f_{\mathrm{c}}$, they report that the distribution stabilizes on [^10] The result was first derived by 2023ApJ...945...86L in the context of stellar EMRIs. $P(e_{\mathrm{pl}})=2g(e_{\mathrm{pl}})g^{\prime}(e_{\mathrm{pl}})$, now being independent on the value of $f_{\bullet}$, where
@@
and $g^{\prime}(e)$ is its derivative with respect to $e$. Here they make use of [1964PhRv..136.1224P]
@@
Although the general trends of their distribution are qualitatively comparable to ours, they feature a peak at smaller $e_{\mathrm{pl}}$ and a slower decay as $e_{\mathrm{pl}}\rightarrow 1$ compared to what we observe. Even for $M_{\bullet}=10^{4}\,\mathrm{M_{\odot}}$, for which the location of the peaks is compatible, our decay at large eccentricity is better fitted by an exponential as opposed to a power law. The incompatibility is also caused by the fact that their $P(e_{\mathrm{pl}})$ does not smoothly go to 0 when $e_{\mathrm{pl}}\to 1$, and instead is sharply cut at some maximum value. For example, their model predicts that (independently on the MBH mass) $1\%$ of EMRIs plunge at $e_{\mathrm{pl}}>0.9$ if $f_{\bullet}=10^{-4}$, and $7\%$ do the same if $f_{\bullet}>f_{\mathrm{c}}$. For comparison, we get between $0.1\%$ and $0.5\%$ depending on $M_{\bullet}$ for the same computation.